### v4

@@ Introduction
The landscape of scientific publishing has been profoundly reshaped by the rise of preprint servers, which facilitate rapid dissemination of research findings across numerous platforms (1, 2, 3) (Fig. Methods, Fig. Methods). This acceleration, however, has transferred the complex task of typesetting from publishers to researchers (4, 5, 6). For those of us in computational fields, this is compounded by a more pressing challenge: ensuring the manuscript remains perfectly synchronised with our data and analysis code. We have all faced the tedious and error-prone task of manually updating a p-value or a sample size in the text after re-running an analysis (7). This disconnect between the research and the report is a critical issue, particularly in disciplines like bioimage analysis, where our findings are built upon complex computational pipelines (8, 9).
Figure 1: System Architecture. Rxiv-Maker integrates Markdown content, YAML metadata, executable scripts, and bibliographies through a processing engine that combines local execution with LaTeX compilation to produce publication-ready PDFs.
Figure 2: Processing Pipeline. User-provided content (left) undergoes automated processing (right), including parsing, script execution, LaTeX conversion, and PDF compilation.
To address these pain points, we developed Rxiv-Maker, a framework designed to streamline the creation of scientific manuscripts. It automates the production of publication-ready PDFs from simple Markdown, harnessing the typesetting power of LaTeX without requiring the author to write LaTeX code. Rxiv-Maker transforms the document into a self-updating manuscript by executing embedded Python or R scripts during compilation. This approach forges a direct, traceable link between the data and the final publication, eliminating manual transcription errors and ensuring the document is always synchronised with the latest results. The framework is designed as a local-first system, providing a rapid and responsive authoring experience that integrates seamlessly with version control systems like Git, thereby enhancing transparency and collaboration (10, 11). The overall architecture and processing pipeline are illustrated in Fig. 1 and Fig. 2, respectively. A tutorial is provided to guide new users through their first manuscript (Methods).
@@ 0.0.1 From Static Text to Self-Updating Manuscripts
A central achievement of Rxiv-Maker is its ability to create dynamic, self-updating documents. By embedding executable code snippets directly within the Markdown source, the manuscript transitions from a static report to a living document. This functionality enables the direct execution of analysis scripts during compilation, with the results (whether statistical values, tables, or figures) directly injected into the text. This process eliminates the possibility of transcription errors when, for example, copying a value from a terminal or another program into the manuscript. We demonstrate this with a practical example from this very paper. The Python code shown in Fig. 3 is embedded in the source of our manuscript. At build time, it retrieves the latest public data on arXiv submissions, calculates key statistics, and inserts them into the text. This ensures that our discussion of preprint growth is always based on the most current evidence available. The analysis reveals that, since 1991, a total of 2.84 million submissions have been made over 35 years (Fig. Methods). These are not static numbers; they were computed when this document was compiled on September 26, 2025, using data updated in September 2025. This capability extends to all forms of programmatic figure generation, as detailed in Methods, ensuring that every visual and numerical claim is reproducibly generated.
@@
0.0.2 Automated Professional Typesetting from Simple Text
Rxiv-Maker allows researchers to achieve the typographic quality of LaTeX while writing in simple, intuitive Markdown. This is accomplished through a sophisticated, multi-pass translator that converts Markdown to LaTeX in a guarded manner. We recognised that a naive, single-pass conversion would be too fragile for academic documents, where complex syntax for mathematics, citations, and cross-references must be preserved. Our translator first identifies and protects delicate elements: mathematical expressions (Methods), code blocks, and citation keys. Subsequent passes then normalise document structure and intelligently convert Rxiv-Maker’s extended Markdown syntax into the corresponding LaTeX commands, supporting comprehensive text formatting including bold, italic, underlined, and other typographic elements. The system handles complex nested formatting combinations such as bold within underline and underlined within bold seamlessly. For instance, it transforms image syntax into floating figure environments, automatically handling captions, labels, and layout parameters. This process enables authors to specify figure widths or positions using simple options in Markdown, without needing to modify LaTeX code. The system supports a wide array of figure generation methods, from static images to script-based visualisations (Table Methods), ensuring all visual elements are correctly and professionally typeset. For more advanced formatting needs, including complex tables, users can directly inject LaTeX code using {{tex:...}} blocks. A complete syntax reference and example are detailed in Methods.
@@ Discussion
Rxiv-Maker finds its place in a growing ecosystem of tools designed to enhance scientific publishing. Our approach is distinct from collaborative web-based editors like Overleaf (12), which, while excellent for team-based LaTeX writing, do not address the fundamental challenge of keeping a manuscript synchronised with external data and analysis code (13). Our philosophy is more closely aligned with that of platforms like Manubot (14), Jupyter Book (15), and Quarto (16), which also integrate code and narrative, often building on notebook-based formats (17). However, Rxiv-Maker is specifically optimised for the production of high-quality, submission-ready manuscripts from a local-first, developer-centric workflow. Beyond preprint distribution, rxiv-maker excels at generating professional manuscripts suitable for first submission to peer-reviewed journals, providing publication-quality PDFs and clean LaTeX source code that many journals readily accept prior to their final journal-specific formatting requirements. This positions rxiv-maker as a comprehensive solution spanning the entire academic publishing pipeline—from initial draft through preprint distribution to journal submission. As shown in our comparison with other tools (Table Methods), its unique strengths lie in its self-updating capabilities, its intelligent caching for rapid rebuilds, and its deep integration with the Git-based workflows that are common in computational research. We also note other modern typesetting systems like Typst (18) and frameworks like Bookdown (19), which offer alternative paradigms for scientific document creation. While powerful, Rxiv-Maker has limitations. Its primary output is currently PDF, which is ideal for preprints but less suited for the web. To address this, we plan to integrate Pandoc (20) into our pipeline, which will enable multi-format outputs, such as HTML and EPUB, while preserving our commitment to typographic quality. We also intend to enhance the framework’s integration with computational environment managers to further improve the portability and reproducibility of our executable manuscripts across different systems and institutions (8, 9), for which containerisation technologies are also a valid alternative (21, 22). By applying the principles of literate programming (23) to the creation of academic papers, Rxiv-Maker helps bridge the gap between computational research and traditional scientific publishing. It fosters a workflow where the manuscript is not merely a report of the research, but a reproducible and verifiable embodiment of it. In an era where the responsibility for typesetting and validation increasingly falls to authors, tools that automate quality control and promote rigour are essential. Rxiv-Maker is our contribution to this movement, offering a framework that empowers researchers to produce transparent, reproducible, and professionally polished manuscripts with confidence and efficiency.
@@ Acknowledgements.
[ht] Tool Type Markdown Primary Use Case Key Strengths Open Source Rxiv-Maker Pipeline Excellent Reproducible preprints Local-first execution, automated caching, rich CLI Yes Overleaf (12) Web Editor Limited Collaborative LaTeX Real-time collaboration, rich templates, cloud-based Freemium Quarto (16) Publisher Native Multi-format publishing Polyglot support, multiple outputs, scientific focus Yes Manubot (14) Collaborative Native Version-controlled writing Automated citations, transparent collaboration, Git-based Yes Pandoc (20) Converter Excellent Format conversion Universal format support, extensible filters Yes Typst (18) Typesetter Good Modern typesetting Fast compilation, modern syntax, growing ecosystem Yes Bookdown (19) Publisher R Markdown Academic books Cross-references, multiple formats Yes Direct LaTeX Typesetter None Traditional publishing Full control, established workflows, mature ecosystem Yes Comparison of Manuscript Preparation Tools. This comparison positions each tool within the scientific publishing ecosystem. Rxiv-Maker specialises in reproducible preprint workflows with local-first execution and developer-centric features. Other tools address distinct needs, such as real-time collaborative editing (Overleaf), multi-format output (Quarto), or version-controlled writing (Manubot).
@@ A Getting Started Tutorial
This straightforward process allows you to get from a blank slate to a compiled PDF in minutes, providing a solid foundation that you can then build upon with more advanced features like programmatic figures and tables.
@@
For script-based figures, you place your .py or .R scripts in the FIGURES directory. These scripts often leverage powerful plotting libraries such as Matplotlib (24) or Seaborn (25). During compilation, Rxiv-Maker executes these scripts, and any image files they save (e.g., PNG, PDF, SVG) are automatically detected and can be included in your manuscript. This ensures your visualisations are always synchronised with the underlying data and analysis, as a change in one will trigger the regeneration of the other. This is the method used to produce Fig. Methods and Fig. Methods.
For diagrams, such as flowcharts or system architectures, you can use Mermaid (26). You create a .mmd file containing a text-based description of your diagram. The framework uses the Mermaid command-line tool to render this description into a vector or raster image. This allows your diagrams to be version-controlled just like code, making them easy to modify and track over time.
@@
{sfigure}
[p] [Uncaptioned image] The growth of preprint submissions on the arXiv server (1991-2025). This figure was generated from public arXiv statistics using a Python script executed by the Rxiv-Maker pipeline, demonstrating reproducible, data-driven visualisation.
{sfigure}
[p] [Uncaptioned image] Preprint Submission Trends Across Multiple Servers (2018-2025). This figure, showing preprints indexed by PubMed from major repositories, was generated from public data (27) using a reproducible R script within the Rxiv-Maker pipeline.
{sfigure}
[p] [Uncaptioned image] The Rxiv-Maker Visual Studio Code Extension. The extension enhances the authoring experience by providing syntax highlighting for extended Markdown, autocompletion for citation keys (e.g., @Knuth...), and real-time validation to catch errors as you type.
{sfigure}
[p] [Uncaptioned image] Detailed System Architecture. A comprehensive technical diagram of the Rxiv-Maker architecture, illustrating the input layers, the processing engine’s components (parsers, converters, generators), the compilation infrastructure, and the final output generation. This figure highlights the modular design that enables robust and reproducible manuscript compilation.

### v5

@@ Introduction
The landscape of scientific publishing has been profoundly reshaped by the rise of preprint servers, which facilitate rapid dissemination of research findings across numerous platforms beck2020; levchenk2024; Fraser2020_preprint_growth (Fig. Methods, Fig. Methods). This acceleration, however, has transferred the complex task of typesetting from publishers to researchers Vale2015_preprints; Tenant2016_academic_publishing; lin2020. For those of us in computational fields, this is compounded by a more pressing challenge: ensuring the manuscript remains perfectly synchronised with our data and analysis code. We have all faced the tedious and error-prone task of manually updating a p-value or a sample size in the text after re-running an analysis perkel2022. This disconnect between the research and the report is a critical issue, particularly in disciplines like bioimage analysis, where our findings are built upon complex computational pipelines biaflows2024; dl4miceverywhere2024.
Figure 1: System Architecture. Rxiv-Maker integrates Markdown content, YAML metadata, executable scripts, and bibliographies through a processing engine that combines local execution with LaTeX PDF compilation or Word (.docx) generation to produce publication-ready documents.
Figure 2: Processing Pipeline. User-provided content (left) undergoes automated processing (right), including parsing, script execution, document compilation (PDF/Word), and final output generation.
To address these challenges, we developed Rxiv-Maker, a framework designed to streamline the creation of scientific manuscripts. It automates the production of publication-ready documents from simple Markdown, harnessing the typesetting power of LaTeX or the flexibility of Microsoft Word without requiring the author to manage complex formatting manually. Rxiv-Maker transforms the document into a self-updating manuscript by executing embedded Python or R scripts during compilation. This approach forges a direct, traceable link between the data and the final publication, eliminating manual transcription errors and ensuring the document is always synchronised with the latest results. The framework is designed as a local-first system, providing a rapid and responsive authoring experience that integrates seamlessly with version control systems like Git, thereby enhancing transparency and collaboration Ram2013_git_science; Perez-Riverol2016_github_bioinformatics. The overall architecture and processing pipeline are illustrated in Fig. 1 and Fig. 2, respectively. A tutorial is provided to guide new users through their first manuscript (Methods).
@@ 0.0.1 From Static Text to Self-Updating Manuscripts
A central achievement of Rxiv-Maker is its ability to create dynamic, self-updating documents. By embedding executable code snippets directly within the Markdown source, the manuscript transitions from a static report to a living document. This functionality enables the direct execution of analysis scripts during compilation, with the results (whether statistical values, tables, or figures) directly injected into the text. This process eliminates the possibility of transcription errors when, for example, copying a value from a terminal or another program into the manuscript. We demonstrate this with a practical example from this very paper. The Python code shown in Fig. 3 is embedded in the source of our manuscript. At build time, it retrieves the latest public data on arXiv submissions, calculates key statistics, and inserts them into the text. This ensures that our discussion of preprint growth is always based on the most current evidence available. The analysis reveals that, since 1991, a total of 2.91 million submissions have been made over 35 years (Fig. Methods). These are not static numbers; they were computed when this document was compiled on December 17, 2025, using data updated in December 2025. This capability extends to all forms of programmatic figure generation, as detailed in Methods, ensuring that every visual and numerical claim is reproducibly generated.
@@
0.0.2 Multi-Format Production: From PDF to Word
Rxiv-Maker allows researchers to achieve the typographic quality of LaTeX while writing in simple, intuitive Markdown. This is accomplished through a sophisticated translator that converts Markdown to restricted LaTeX in a guarded manner. We recognised that a naive, single-pass conversion would be too fragile for academic documents, where complex syntax for mathematics, citations, and cross-references must be preserved. Our translator first identifies and protects delicate elements: mathematical expressions (Methods), code blocks, and citation keys. Subsequent passes then normalise document structure and intelligently convert Rxiv-Maker’s extended Markdown syntax into the corresponding LaTeX commands, supporting comprehensive text formatting.
Importantly, Rxiv-Maker extends beyond PDF generation to support the collaborative realities of modern science, where co-authors or journals may require editable formats. The framework includes a dedicated Word export engine (built on python-docx) that maps the same semantic Markdown source to .docx documents. This ensures that the generated Word files maintain figure numbering, citations, and equations, allowing for seamless interchange with collaborators. Furthermore, to facilitate the review process, the track-changes feature leverages Git history to generate annotated PDFs (using latexdiff), enabling authors to visually visualize text changes between manuscript versions or against specific milestones (e.g., submission vs. revision).
The system handles complex nested formatting, inline DOI resolution, and transforms image syntax into floating figure environments, automatically handling captions, labels, and layout parameters. This enables authors to specify figure widths or positions using simple options in Markdown, without needing to modify LaTeX code. The system supports a wide array of figure generation methods, from static images to script-based visualisations (Table Methods), ensuring all visual elements are correctly and professionally typeset. For more advanced formatting needs, including complex tables, users can directly inject LaTeX code using {{tex:...}} blocks. A complete syntax reference and example are detailed in Methods.
@@ Discussion
Rxiv-Maker finds its place in a growing ecosystem of tools designed to enhance scientific publishing. Our approach is distinct from collaborative web-based editors like Overleaf Overleaf2024, which, while excellent for team-based LaTeX writing, do not address the fundamental challenge of keeping a manuscript synchronised with external data and analysis code HenriquesLab2015_template. Our philosophy is more closely aligned with that of platforms like Manubot himmelstein2019, Jupyter Book JupyterBook2020, and Quarto Quarto2024, which also integrate code and narrative, often building on notebook-based formats Jupyter2016_notebook. However, Rxiv-Maker is specifically optimised for the production of high-quality, submission-ready manuscripts from a local-first, developer-centric workflow. Beyond preprint distribution, rxiv-maker excels at generating professional manuscripts suitable for first submission to peer-reviewed journals, providing publication-quality PDFs and clean LaTeX source code that many journals readily accept prior to their final journal-specific formatting requirements. This positions rxiv-maker as a comprehensive solution spanning the entire academic publishing pipeline—from initial draft through preprint distribution to journal submission. As shown in our comparison with other tools (Table Methods), its unique strengths lie in its self-updating capabilities, its intelligent caching for rapid rebuilds, and its deep integration with the Git-based workflows that are common in computational research. We also note other modern typesetting systems like Typst Typst2024 and frameworks like Bookdown Xie2016_bookdown, which offer alternative paradigms for scientific document creation. While powerful, Rxiv-Maker continues to evolve. Recent updates have addressed the need for broader collaboration by introducing Word export capabilities, allowing authors to share manuscripts with collaborators who rely on word processors. We also plan to integrate Pandoc pandoc2020 more deeply to enable further multi-format outputs, such as HTML and EPUB, while preserving our commitment to typographic quality. We also intend to enhance the framework’s integration with computational environment managers to further improve the portability and reproducibility of our executable manuscripts across different systems biaflows2024; dl4miceverywhere2024, for which containerisation technologies are also a valid alternative Boettiger2015_docker_reproducibility; gomes2018. By applying the principles of literate programming Knuth1984_literate_programming to the creation of academic papers, Rxiv-Maker helps bridge the gap between computational research and traditional scientific publishing. It fosters a workflow where the manuscript is not merely a report of the research, but a reproducible and verifiable embodiment of it. In an era where the responsibility for typesetting and validation increasingly falls to authors, tools that automate quality control and promote rigour are essential. Rxiv-Maker is our contribution to this movement, offering a framework that empowers researchers to produce transparent, reproducible, and professionally polished manuscripts with confidence and efficiency.
@@ Acknowledgements.
[ht] Tool Type Markdown Primary Use Case Key Strengths Open Source Rxiv-Maker Pipeline Excellent Reproducible preprints Local-first execution, automated caching, rich CLI Yes Overleaf Overleaf2024 Web Editor Limited Collaborative LaTeX Real-time collaboration, rich templates, cloud-based Freemium Quarto Quarto2024 Publisher Native Multi-format publishing Polyglot support, multiple outputs, scientific focus Yes Manubot himmelstein2019 Collaborative Native Version-controlled writing Automated citations, transparent collaboration, Git-based Yes Pandoc pandoc2020 Converter Excellent Format conversion Universal format support, extensible filters Yes Typst Typst2024 Typesetter Good Modern typesetting Fast compilation, modern syntax, growing ecosystem Yes Bookdown Xie2016_bookdown Publisher R Markdown Academic books Cross-references, multiple formats Yes Direct LaTeX Typesetter None Traditional publishing Full control, established workflows, mature ecosystem Yes Comparison of Manuscript Preparation Tools. This comparison positions each tool within the scientific publishing ecosystem. Rxiv-Maker specialises in reproducible preprint workflows with local-first execution and developer-centric features. Other tools address distinct needs, such as real-time collaborative editing (Overleaf), multi-format output (Quarto), or version-controlled writing (Manubot).
@@ A Getting Started Tutorial
This straightforward process allows you to get from a blank slate to a compiled PDF, providing a solid foundation that you can then build upon with more advanced features like programmatic figures and tables.
@@
For script-based figures, you place your .py or .R scripts in the FIGURES directory. These scripts often leverage powerful plotting libraries such as Matplotlib Hunter2007_matplotlib or Seaborn Waskom2021_seaborn. During compilation, Rxiv-Maker executes these scripts, and any image files they save (e.g., PNG, PDF, SVG) are automatically detected and can be included in your manuscript. This ensures your visualisations are always synchronised with the underlying data and analysis, as a change in one will trigger the regeneration of the other. This is the method used to produce Fig. Methods and Fig. Methods.
For diagrams, such as flowcharts or system architectures, you can use Mermaid Mermaid2023_documentation. You create a .mmd file containing a text-based description of your diagram. The framework uses the Mermaid command-line tool to render this description into a vector or raster image. This allows your diagrams to be version-controlled just like code, making them easy to modify and track over time.
@@
{sfigure*}
[p!] [Uncaptioned image] The growth of preprint submissions on the arXiv server (1991-2025). This figure was generated from public arXiv statistics using a Python script executed by the Rxiv-Maker pipeline, demonstrating reproducible, data-driven visualisation.
{sfigure*}
[p!] [Uncaptioned image] Preprint Submission Trends Across Multiple Servers (2018-2025). This figure, showing preprints indexed by PubMed from major repositories, was generated from public data PubMedByYear2025 using a reproducible R script within the Rxiv-Maker pipeline.
{sfigure*}
[p!] [Uncaptioned image] The Rxiv-Maker Visual Studio Code Extension. The extension enhances the authoring experience by providing syntax highlighting for extended Markdown, autocompletion for citation keys (e.g., @Knuth...), and real-time validation to catch errors as you type.